Hi,

We did linear equations last week in class, and now we've moved on to graphing, but I can't do my homework because I still don't understand the linear equations yet!

On the blackboard, our teacher explained that the answer to:

x + 2y = 0

is

y = -1/2x

If someone could explain how our teacher got to that answer, I'd REALLY appreciate it!!

Also, the book we're using said to just guess a number for x, and then figure out y... does that mean everyone will get a different answer?

Thanks,
~ Jasmine

Answer by checkley71(8403)

X+2Y=0 OR 2Y=-X+0 OR Y=-X/2 OR Y=-(1/2)X
IF YOU PICK A VALUE FOR Y AND SOLVE FOR X THIS WILL GIVE YOU A POINT ON THE LINE. THERE ARE MANY POINTS ON THE LINE THEREFORE MANY DIFFERENT VALUES FOR X7Y
EXAMPLE LET Y=10 THEN 10=-(1/2)X OR OR DIVIDING BOTH SIDES BY -1/2 WE GET
10/-.5=X OR -20=X THUS THIS POINT IS (-20,10)
LET X=4 AND WE GET Y=-(1/2)4 OR Y=-.5*4 OR Y=-2 THUS THIS POINT IS (4,-2)
HOPE THIS HELPS